Mrs. Brook lacks 7 years from being 5 times as old as her son. Six years from now she will lack 3 years from being three times as old as her son is then. Find each of their ages.
Answer by Starz(4)

let's see...we can work this out with an equation
let x represent the age of her son
5x-7=3x-3+6
now....let's solve
5x-7=3x-3+6
5x=3x-3+6+7
5x=3x+16
5x-3x=16
2x=16
x=16/2
x=8
Now that you have the son's age, you can find out mrs. brook's age.
5 x 8-7
40-7
33
Therefore, her son is 8 and she is 33.